# EDiT: Efficient Diffusion Transformers with Linear Compressed Attention

### v1

@@ Abstract
Diffusion Transformers (DiTs) have emerged as a leading architecture for text-to-image synthesis, producing high-quality and photorealistic images. However, the quadratic scaling properties of the attention in DiTs hinder image generation with higher resolution or on devices with limited resources. This work introduces an efficient diffusion transformer (EDiT) to alleviate these efficiency bottlenecks in conventional DiTs and Multimodal DiTs (MM-DiTs). First, we present a novel linear compressed attention method that uses a multi-layer convolutional network to modulate queries with local information while keys and values are spatially aggregated. Second, we formulate a hybrid attention scheme for multi-modal inputs that combines linear attention for image-to-image interactions and standard scaled dot-product attention for interactions involving prompts. Merging these two approaches leads to an expressive, linear-time Multimodal Efficient Diffusion Transformer (MM-EDiT). We demonstrate the effectiveness of the EDiT and MM-EDiT architectures by integrating them into PixArt-$\Sigma$ (conventional DiT) and Stable Diffusion 3.5-Medium (MM-DiT), achieving up to $\times2$ speedup with comparable image quality after distillation.
@@ 1 Introduction
Diffusion Transformers (DiTs) [16], have quickly become the de-facto choice for architecture and are prevalent within state-of-the-art (SOTA) text-to-image generation models such as SANA [22], FLUX [11], PixArt-$\Sigma$ [2], Stable Diffusion v3 [5], and Lumina-Next [24]. The majority of these models rely on standard quadratic scaled dot-product attention [20] to capture long-range dependencies and global context, enabling them to generate high-quality, photorealistic, and aesthetically pleasing images. Earlier versions of DiTs relied mainly on the conventional transformer architecture, i.e., image tokens were processed through self-attention, whereas prompt conditioning was achieved by applying cross-attention [3, 2]. More recently, Multimodal DiTs (MM-DiTs) have introduced joint attention mechanisms to handle both image and prompt tokens within a single self-attention layer [5]. While enhancing image quality significantly, joint attention scales poorly due to the increased number of involved tokens. This limitation is further aggravated when generating high-resolution images, mainly due to the increase in the amount of input tokens, prohibiting the efficient deployment of SOTA text-to-image models, particularly on resource-constrained devices.
Existing research explored strategies to alleviate this computational bottleneck within DiTs primarily by focusing on either linearizing the attention mechanism [22, 14] or reducing the number of tokens [2]. For instance, Linfusion [14] relies on multi-layer non-linear transformations to map input tokens, thereby enhancing the expressiveness of the linear attention mechansim. Whereas, SANA [22] combines standard linear attention [10] with convolutional operations in the feed forward part of the transformer, leveraging the inherent structure of images while capturing local relationships more effectively. Recent work, e.g., PixArt-$\Sigma$ [2], reduces the number of keys and values involved in the computations, which is an effective strategy to improve computational efficiency.
@@
To efficiently handle multimodal inputs, we further introduce a hybrid attention mechanism that combines the strengths of linear and scaled dot-product attention to handle MM-DiT based models. In particular, we include EDiT’s linear compressed attention for the computationally intensive interactions, i.e., when computing image-to-image dependencies, on the other hand, scaled dot-product attention is used for capturing text interactions. These modifications introduce a novel Multimodal Efficient Diffusion Transformer (MM-EDiT) architecture, which achieves linear time complexity with respect to image size without sacrificing quality. We demonstrate the effectiveness of our hybrid attention approach by distilling Stable Diffusion 3.5-Medium, an MM-DiT-based model, and show that MM-EDiT outperforms existing linear attention alternatives. Our contributions can be summarized as follows:
@@
    We propose MM-EDiT, integrating our linear compressed attention into MM-DiTs using a novel hybrid attention formulation (Section 3.2).
@@
    Lastly, we demonstrate the effectiveness of these novel architectures by adapting PixArt-$\Sigma$ [2] and StableDiffusion 3.5-Medium through distillation (Section 4). Our extensive experiments show that EDiT and MM-EDiT can yield a speedup of up to $\times2$ without compromising on image quality.
@@ 2 Background
While this mechanism is highly successful, it requires the explicit computation of $QK^{T}$, which scales quadratically with the number of tokens and presents a major computational bottleneck in self-attention for approaches using scaled dot-product attention. To address this issue, several alternative attention mechanisms [10, 7] have been proposed that scale linearly with the number of tokens rather than quadratically.
@@ 2.1 Linear Attention
$Y=\{y_{i}\}_{i=1:N}~{}~{}\text{where}~{}~{}y_{i}=\frac{\sum_{j=1}^{N}\operatorname{sim}\left(Q_{i},K_{j}\right)V_{j}}{\sum_{j=1}^{N}\operatorname{sim}\left(Q_{i},K_{j}\right)}.$ (2)
If we substitute the similarity function $\operatorname{sim}(q,k)=\exp\left(\frac{q^{T}k}{\sqrt{d}}\right)$, Equation (2) is equivalent to the scaled dot-product attention in Equation (1). Generalizing this even further, self-attention can be expressed as a linear dot product of kernel feature maps: $\exp\left(\frac{q_{i}\cdot k_{j}}{\sqrt{d}}\right)v_{j}\approx(\phi\left(q_{i}\right)\cdot\phi\left(k_{j}\right))\cdot v_{j}$, which can be further simplified to $\phi\left(q_{i}\right)\cdot(\phi\left(k_{j}\right)\cdot v_{j})$ by applying the associativity property of matrix multiplication. In most linear attention works, different kernel functions are considered to formulate $K=\phi_{K}(X)$, $Q=\phi_{Q}(X)$, and $V=\phi_{V}(X)$. The resulting formula for linear attention is expressed as:
@@
where, $y_{i}$ denotes a single token in the output sequence $Y$. While the formulation of linear attention in Equation (3) remains consistent across studies, the literature offers a variety of approaches to define the kernel feature maps $\phi_{Q}$, $\phi_{K}$, and $\phi_{V}$. Some works have demonstrated that using basic formulations for $\phi_{K}$ and $\phi_{Q}$ with functions such as $\textrm{ReLU}(x)$ or $\textrm{ELU}(x)+1$ while keeping $\phi_{V}$ linear can yield performance comparable to scaled dot-product attention.
@@ 2.2 Efficient Attention for Diffusion Models
LinFusion [14] explores linear attention for UNet-based diffusion approaches using Stable Diffusion v1.5 [18] and Stable Diffusion XL [17]. To this end, Liu et al. compute queries and keys with the following function:
$\phi_{\textrm{LF}}(x_{n})=1+$
$\qquad\,\,\textrm{ELU}(x_{n}+\textrm{Linear}(\textrm{LN}(\textrm{LeakyReLU}(\textrm{Linear}(x_{n}))))),$
where LN denotes layer normalization [1]. The first linear layer projects each $x_{n}$ to a lower dimensional space, while the second linear layer projects back to the original space. For the values, they use a linear projection. This linear attention mechanism is then integrated into existing models, which are subsequently distilled. Although, this approach works well for UNets, results presented in both  [14] and in this paper (see Table 1) show that it fails for DiTs.
Key–Value Token Compression. Another approach to accelerate attention is key-value (KV) token compression, which reduces the number of keys and value tokens. This technique is implemented in the DiT-based PixArt-$\Sigma$ [2]. Chen et al. use a depthwise convolution to compress keys and values. In practice, this operation aggregates $k\times k$ blocks of tokens into a single compressed token, which lowers the amount of keys and values by a factor of $k$. Hence, the size of the matrix $QK^{T}$ is also smaller by a factor of $k$, thus improving efficiency. In PixArt-$\Sigma$, the same compression kernel is used for both keys and values. As this kernel is integrated with conventional self-attention, the computational complexity remains quadratic.
Figure 2: (a) Scaled Dot Product Joint Attention: Image-to-Image attention matrix $(Q_{I}K_{I}^{T})$ grows quadratically with number of image tokens. (b) MM-EDiT Attention: Our hybrid attention approach combines scaled dot product attention and linear attention, which reduces the time complexity of attention to $\mathcal{O}(N_{I})$. We use ConvFusion for $Q_{I}$, while $K_{I}$ and $V_{I}$ are obtained through Spatial Compressor.
@@ 3.1 EDiT - Linear Attention for DiTs
EDiT’s linear compressed attention processes queries with multi-layered convolutions and compresses key and value tokens spatially using convolution-based feature maps.
Processing Queries. Inspired by the intuition behind LinFusion [14] and SANA [22], we propose the ConvFusion function. This operation effectively combines LinFusion and SANA’s MixFFN into a unified convolutional mapping from tokens to queries. In particular, we process queries using $\phi_{\textrm{CF}}(\cdot)$ from input tokens $X$ using:
$Q^{\textrm{EDiT}}=\phi_{\textrm{CF}}(X)=$
$\textrm{ReLU}(X+\textrm{Conv}(\textrm{GN}(\textrm{LeakyRELU}(\textrm{Conv}(X))))),$ (4)
@@
[…]
@@ 3.2 MM-EDiT - Hybrid Attention for MM-DiTs
where we define $\eta_{I}=\eta(Q_{I},K_{I},K_{P})$ and $\eta_{P}=\eta(Q_{P},K_{P},K_{I})$. With this decomposition, the softmax in $\mathbf{A}(\cdot,\cdot,\cdot)$ is now applied over different token sets compared to joint attention, resulting in row sums that are not equal to 1. Therefore, a normalization factor, $\eta(\cdot,\cdot,\cdot)$, as shown in Equation (7), needs to be added to adjust each element of the attention matrix so that the output remains unchanged:
$\eta(\mathbf{q},\mathbf{k}_{1},\mathbf{k}_{2})=\dfrac{\sum_{j}\exp(\mathbf{q}\mathbf{k}_{1}^{T})_{j}}{\sum_{j}\exp(\mathbf{q}\mathbf{k}_{1}^{T})_{j}+\sum_{j^{\prime}}\exp(\mathbf{q}\mathbf{k}_{2}^{T})_{j^{\prime}}}.$ (7)
@@
We begin by computing $Q_{I}$ using ConvFusion (Equation (3.1)) and obtaining $K_{I}$ and $V_{I}$ through Spatial Compressor (Equation (5)). For prompt tokens, $Q_{P}$, $K_{P}$, and $V_{P}$ are derived using simple linear projections, as in the original model. Following the MM-DiT approach, we then concatenate the image and prompt tokens. Finally, we replace $\mathbf{A}(Q_{I},K_{I},V_{I})$ in the scaled dot-product joint attention (Equation (3.2)) with $\mathbf{A}^{\textrm{Lin}}(Q_{I},K_{I},V_{I})$, resulting in the hybrid joint attention matrix $\mathbf{A}^{\textrm{Hybrid}}(Q,K,V)$ formulated as follows:
@@ 4 Experiments
We evaluate our EDiT and MM-EDiT models by distilling two well-known diffusion transformers using our novel EDiT and MM-EDiT architectures. First, we use PixArt-$\Sigma$ [2] as a conventional DiT, and second, we use StableDiffusion 3.5-Medium (SD-v3.5M) [5] as the basis for our MM-DiT experiments. Dataset. In all experiments, we rely on the YE-POP [^1] https://huggingface.co/datasets/Ejafa/ye-pop dataset for training. This subset of Laion-POP comprises about 480K images with two alternative captions each. For evaluation, we use the PixartEval30K dataset [2] [^2] https://huggingface.co/datasets/PixArt-alpha/PixArt-Eval-30K.
CLIP $\uparrow$ FID $\downarrow$ FID $\downarrow$ CLIP $\uparrow$ FID $\downarrow$ FID $\downarrow$ Latency (s) $\downarrow$
@@
PixArt-$\Sigma$ 0.285 07.57 02.50 0.285 07.09 02.53 05.44
EDiT (ours) 0.283 07.06 02.57 0.290 07.82 02.64 04.15
SANA-DiT 0.283 08.43 03.31 0.286 09.31 03.16 04.00
LinFusion-DiT 0.289 15.87 05.98 0.283 44.66 11.01 04.20
KV Comp. ($k=2$) 0.275 10.69 03.77 0.283 10.32 03.50 05.00
@@
0.284 07.59 02.59 0.286 07.76 02.72 04.55
0.286 14.53 04.82 0.290 26.59 06.20 03.80
0.284 07.06 02.53 0.287 07.70 02.69 04.30
Table 1: Quantitative comparison of EDiT to it’s teacher (PixArt-$\Sigma$), prior efficient DiT architectures and alternative design choices at resolutions $512\times 512$ and $1024\times 1024$. We ablate different choices to compute queries, keys and values, where we denote ConvFusion and Spatial Compressor by $CF$ and $SC$. The latency indicates the time to generate a single $1024\times 1024$ pixel image on a consumer grade Nvidia 3090 RTX with $20$ diffusion steps. The results show that EDiT performs similar to the teacher, outperforms previous approaches, and strikes the best compromise of runtime and quality across all considered variants.
Training details for EDiT. We use three publicly available checkpoints for PixArt-$\Sigma$, trained on resolutions $512\times 512$, $1024\times 1024$, and $2048\times 2048$ pixels. We initialize our approach using the $512\times 512$ PixArt-$\Sigma$ checkpoint, modify the self-attention layers using our linear compressed attention approach, and distill the model for 30 epochs. In particular, we follow [14, 12] and combine a noise prediction loss with knowledge, and feature distillation. For the feature level distillation, we use the output of each self-attention layer and minimize the divergence between teacher and student. Finally, as knowledge distillation, we minimize the difference betweent the student and teacher’s prediction.
We take the resulting model, unfreeze all weights of EDiT, and train for 10 more epochs at a resolution of $1024\times 1024$ using the teacher checkpoint for the respective resolution. Finally, we repeat the same process with a resolution of $2048\times 2048$ for $8$ more epochs.
Training details for MM-EDiT. We use Stable Diffusion 3.5 Medium (SD-v3.5M) to evaluate the efficacy of our MM-EDiT mechanism. In contrast to PixArt-$\Sigma$, there is only one publicly available checkpoint for SD-v3.5M, which can be used for multiple resolutions. We take this pretrained SD-v3.5M checkpoint and replace the multimodal attention layers with MM-EDiT. We first train the hybrid attention layers for 45 epochs on $512\times 512$ images using a combined loss similar to that used for EDiT. We again combine a task loss, knowledge, and feature distillation, but this time, the task loss is based on the rectified flow mechanism [15] taken from SD-v3.5M. Following this, we unfreeze all weights of the MM-EDiT modules and train them for 10 more epochs at a resolution of $1024\times 1024$. At this stage, the combined loss showed suoptimal performance, and we switched to pure knowledge distillation.

### v2

@@ Abstract
Diffusion Transformers (DiTs) have emerged as a leading architecture for text-to-image synthesis, producing high-quality and photorealistic images. However, the quadratic scaling properties of the attention in DiTs hinder image generation with higher resolution or on devices with limited resources. This work introduces an efficient diffusion transformer (EDiT) to alleviate these efficiency bottlenecks in conventional DiTs and Multimodal DiTs (MM-DiTs). First, we present a novel linear compressed attention method that uses a multi-layer convolutional network to modulate queries with local information while keys and values are aggregated spatially. Second, we formulate a hybrid attention scheme for multimodal inputs that combines linear attention for image-to-image interactions and standard scaled dot-product attention for interactions involving prompts. Merging these two approaches leads to an expressive, linear-time Multimodal Efficient Diffusion Transformer (MM-EDiT). We demonstrate the effectiveness of the EDiT and MM-EDiT architectures by integrating them into PixArt-$\Sigma$ (conventional DiT) and Stable Diffusion 3.5-Medium (MM-DiT), achieving up to $2.2\times$ speedup with comparable image quality after distillation.
@@ 1 Introduction
Diffusion Transformers (DiTs) [16], have quickly become the de-facto choice of architecture for state-of-the-art (SOTA) text-to-image generation models such as FLUX [11], PixArt-$\Sigma$ [2], Stable Diffusion v3 [5], and Lumina-Next [24]. The majority of these models rely on standard quadratic scaled dot-product attention [20] to capture long-range dependencies and global context, enabling them to generate high-quality, photorealistic, and aesthetically pleasing images. Earlier versions of DiTs relied mainly on the conventional transformer architecture, i.e., image tokens were processed through self-attention, whereas prompt conditioning was achieved by applying cross-attention [3, 2]. More recently, Multimodal DiTs (MM-DiTs) have introduced joint attention mechanisms to handle both image and prompt tokens within a single self-attention layer [5]. While enhancing image quality significantly, joint attention scales poorly due to the increased number of involved tokens. This limitation is further aggravated when generating high-resolution images, mainly due to the increase in the amount of input tokens, prohibiting the efficient deployment of SOTA text-to-image models, particularly on resource-constrained devices.
Existing research has explored strategies to alleviate this computational bottleneck within DiTs primarily by focusing on linearizing the attention mechanism [22, 14] or reducing the number of tokens [2]. For instance, Linfusion [14] relies on multi-layer non-linear transformations to map input tokens, thereby enhancing the expressiveness of the linear attention mechansim. Whereas, SANA [22] combines standard linear attention [10] with convolutional operations in the feed forward part of the transformer, leveraging the inherent structure of images while capturing local relationships more effectively. Recent work, e.g., PixArt-$\Sigma$ [2], reduces the number of key and value tokens to improve computational efficiency.
@@
To efficiently handle multimodal inputs, we introduce a hybrid attention mechanism that combines the strengths of linear and scaled dot-product attention, enabling effective MM-DiT-based models. In particular, we include EDiT’s linear compressed attention for the computationally intensive interactions, i.e., when computing image-to-image dependencies, on the other hand, scaled dot-product attention is used for capturing text interactions. These modifications introduce a novel Multimodal Efficient Diffusion Transformer (MM-EDiT) architecture that achieves linear-time complexity with respect to image size without sacrificing quality. We demonstrate the effectiveness of our hybrid attention approach by distilling Stable Diffusion 3.5-Medium, an MM-DiT-based model, and show that MM-EDiT outperforms existing linear attention alternatives. Our contributions can be summarized as follows:
@@
    We propose $M$-EDiT, integrating our linear compressed attention into $M$-DiTs using a novel hybrid attention formulation (Section 3.2).
@@
    Lastly, we demonstrate the effectiveness of these novel architectures by adapting PixArt-$\Sigma$ [2] and StableDiffusion 3.5-Medium through distillation (Section 4). Our extensive experiments show that EDiT and MM-EDiT can yield a speedup of up to $2.2\times$ without compromising on image quality.
@@ 2 Background
While this mechanism is highly successful, it requires the explicit computation of $QK^{T}$, which scales quadratically with the number of tokens and presents a major computational bottleneck in self-attention for approaches using scaled dot-product attention. To address this issue, several alternative attention mechanisms [10, 7] have been proposed that scale linearly with the number of tokens.
@@ 2.1 Linear Attention
$Y=\{y_{i}\}_{i=1:N}~~\text{where}~~y_{i}=\frac{\sum_{j=1}^{N}\operatorname{sim}\left(Q_{i},K_{j}\right)V_{j}}{\sum_{j=1}^{N}\operatorname{sim}\left(Q_{i},K_{j}\right)}$ (2)
If we substitute the similarity function $\operatorname{sim}(q,k)=\exp\left(\frac{q^{T}k}{\sqrt{d}}\right)$, Equation (2) is equivalent to the scaled dot-product attention in Equation (1). While generalizing this even further, the self-attention can be expressed as a linear dot product of kernel feature maps: $\exp\left(\frac{q_{i}\cdot k_{j}}{\sqrt{d}}\right)v_{j}\approx(\phi\left(q_{i}\right)\cdot\phi\left(k_{j}\right))\cdot v_{j}$, which can be further simplified to $\phi\left(q_{i}\right)\cdot(\phi\left(k_{j}\right)\cdot v_{j})$ by applying the associativity property of matrix multiplication. In most linear attention works, different kernel functions are considered to formulate $K=\phi_{K}(X)$, $Q=\phi_{Q}(X)$, and $V=\phi_{V}(X)$. The resulting formula for linear attention is expressed as:
@@
where $y_{i}$ denotes a single token in the output sequence $Y$. While the formulation of linear attention in Equation (3) remains consistent across studies, the literature offers a variety of approaches to define the kernel feature maps $\phi_{Q}$, $\phi_{K}$, and $\phi_{V}$. Some works have demonstrated that using basic formulations for $\phi_{K}$ and $\phi_{Q}$ with functions such as $\textrm{ReLU}(x)$ or $\textrm{ELU}(x)+1$ while keeping $\phi_{V}$ linear can yield performance comparable to scaled dot-product attention.
@@ 2.2 Efficient Attention for Diffusion Models
LinFusion [14] explores linear attention for UNet-based diffusion approaches using Stable Diffusion v1.5 [18] and Stable Diffusion XL [17]. To this end, Liu et al. compute queries and keys with the following function $\phi_{\textrm{LF}}(x_{n})=$
$1+\textrm{ELU}(x_{n}+\textrm{Linear}(\textrm{LN}(\textrm{LeakyReLU}(\textrm{Linear}(x_{n}))))),$
where LN denotes layer normalization [1]. The first linear layer projects each $x_{n}$ to a lower dimensional space, while the second linear layer projects back to the original space. For the values, they use a linear projection. This linear attention mechanism is then integrated into existing models, which are subsequently distilled. Although, this approach works for UNets, results presented in both  [14] and in this paper (see Table 1) show that it fails for DiTs.
Key–Value Token Compression. Another approach to accelerate attention is key-value (KV) token compression, which reduces the number of keys and value tokens. This technique is implemented in the DiT-based PixArt-$\Sigma$, where Chen et al [2] use a depthwise convolution to compress keys and values. In practice, this operation aggregates $k\times k$ blocks of tokens into a single compressed token, which lowers the amount of keys and values by a factor of $k$. Hence, the size of the matrix $QK^{T}$ is also smaller by a factor of $k$, thus improving efficiency. In PixArt-$\Sigma$, the same compression kernel is used for both keys and values. As this kernel is integrated with conventional self-attention, the computational complexity remains quadratic.
@@ 3.1 EDiT - Linear Attention for DiTs
EDiT’s linear compressed attention processes queries with multi-layered convolutions and compresses key and value tokens spatially using convolution-based feature maps. Processing Queries. Inspired by the intuition behind LinFusion [14] and SANA [22], we propose the ConvFusion function. This operation effectively combines LinFusion and SANA’s MixFFN into a unified convolutional mapping from tokens to queries. In particular, we process queries using $\phi_{\textrm{CF}}(\cdot)$ from input tokens $X$ using:
$Q^{\textrm{EDiT}}=\phi_{\textrm{CF}}(X)=$ (4)
$\textrm{ReLU}(\textrm{Linear}(X+\textrm{Conv}(\textrm{LeakyRELU}(\textrm{GN}(\textrm{Conv}(X)))))),$
@@
$\text{where,}~~\phi_{\textrm{SC}}(X)=\textrm{Conv}(\textrm{Linear}(X)).$
@@ 3.2 MM-EDiT - Hybrid Attention for MM-DiTs
Figure 2: (a) Scaled Dot Product Joint Attention: Image-to-Image attention matrix $(Q_{I}K_{I}^{T})$ grows quadratically with number of image tokens. (b) $M$-EDiT Attention: Our hybrid attention approach combines scaled dot product attention and linear attention, which reduces the time complexity of attention to $\mathcal{O}(N_{I})$. We use ConvFusion for $Q_{I}$, while $K_{I}$ and $V_{I}$ are obtained through Spatial Compressor.
@@
where we define $\eta_{I}=\eta(Q_{I},K_{I},K_{P})$ and $\eta_{P}=\eta(Q_{P},K_{P},K_{I})$. With this decomposition, the softmax in $\mathbf{A}(\cdot,\cdot,\cdot)$ is now applied over different token sets compared to joint attention, resulting in row sums that are not equal to 1. Therefore, a normalization factor, $\eta(\cdot,\cdot,\cdot)$, as shown in Equation (3.2), needs to be added to adjust each element of the attention matrix so that the output remains unchanged:
$\eta(\mathbf{q},\mathbf{k}_{1},\mathbf{k}_{2})=\dfrac{\sum_{j}\exp(\mathbf{q}\mathbf{k}_{1}^{T})_{j}}{\sum_{j}\exp(\mathbf{q}\mathbf{k}_{1}^{T})_{j}+\sum_{j^{\prime}}\exp(\mathbf{q}\mathbf{k}_{2}^{T})_{j^{\prime}}}$
@@
We begin by computing $Q_{I}$ using ConvFusion (Equation (4)) and obtaining $K_{I}$ and $V_{I}$ through Spatial Compressor (Equation (5)). For prompt tokens, $Q_{P}$, $K_{P}$, and $V_{P}$ are derived using simple linear projections, as in the original model. Following the MM-DiT approach, we then concatenate the image and prompt tokens. Finally, we replace $\mathbf{A}(Q_{I},K_{I},V_{I})$ in the scaled dot-product joint attention (Equation (3.2)) with $\mathbf{A}^{\textrm{Lin}}(Q_{I},K_{I},V_{I})$, resulting in the hybrid joint attention matrix $\mathbf{A}^{\textrm{Hybrid}}(Q,K,V)$ formulated as follows:
@@ 4 Experiments
We evaluate our EDiT and MM-EDiT models by distilling two well-known diffusion transformers using our novel EDiT and MM-EDiT architectures. We use PixArt-$\Sigma$ [2] as a conventional DiT, and StableDiffusion 3.5-Medium (SD-v3.5M) [5] as the basis for our MM-DiT experiments. Dataset. In all experiments, we rely on the YE-POP [^1] https://huggingface.co/datasets/Ejafa/ye-pop dataset for training. This subset of Laion-POP comprises about 480K images with two alternative captions each. For evaluation, we use the PixartEval30K dataset [2] [^2] https://huggingface.co/datasets/PixArt-alpha/PixArt-Eval-30K.
CLIP $\uparrow$ FID $\downarrow$ FID $\downarrow$ CLIP $\uparrow$ FID $\downarrow$ FID $\downarrow$
@@
PixArt-$\Sigma$ 0.285 07.57 02.50 0.285 07.09 02.53
EDiT (ours) 0.283 07.06 02.57 0.290 07.82 02.64
SANA-DiT 0.283 08.43 03.31 0.286 09.31 03.16
LinFusion-DiT 0.289 15.87 05.98 0.283 44.66 11.01
KV Comp. ($k=2$) 0.275 10.69 03.77 0.283 10.32 03.50
@@
0.284 07.59 02.59 0.286 07.76 02.72
0.286 14.53 04.82 0.290 26.59 06.20
0.284 07.06 02.53 0.287 07.70 02.69
Table 1: Quantitative comparison of EDiT to it’s teacher (PixArt-$\Sigma$), prior efficient DiT architectures and alternative design choices at resolutions $512\times 512$ and $1024\times 1024$. We ablate different choices to compute queries, keys and values, where we denote ConvFusion and Spatial Compressor by CFCF and $SC$. The results show that EDiT performs similar to the teacher, outperforms previous approaches.
Training details for EDiT. We use three publicly available checkpoints for PixArt-$\Sigma$, trained on $512,1024$, and $2048$ image resolutions. We initialize our approach using the $512\times 512$ PixArt-$\Sigma$ checkpoint, modify the self-attention layers using our linear compressed attention approach, and distill the model for 30 epochs. In particular, we follow [14, 12] and combine a noise prediction loss with knowledge distillation, where we minimize the difference between the student and teacher’s prediction, and feature distillation with the output of each self-attention between teacher and student. Thereafter, we take the resulting model, unfreeze all weights of EDiT, and train for 10 more epochs at $1024$ resolution using the teacher checkpoint for the respective resolution. We repeat this process with $2048$ resolution images for $8$ more epochs.
Training details for MM-EDiT. We use Stable Diffusion 3.5 Medium (SD-v3.5M) to evaluate the efficacy of our MM-EDiT mechanism. In contrast to PixArt-$\Sigma$, there is only one publicly available checkpoint for SD-v3.5M, which can be used for multiple resolutions. We take this pretrained SD-v3.5M checkpoint and replace the multimodal attention layers with MM-EDiT. We first train the hybrid attention layers for 45 epochs on $512\times 512$ images using a combined loss similar to that used for EDiT. We again combine a task loss, knowledge, and feature distillation, but this time, the task loss is based on the rectified flow mechanism [15] taken from SD-v3.5M. Following this, we unfreeze all weights of the MM-EDiT modules and train them for 10 more epochs at a resolution of $1024\times 1024$ only with knowledge distillation loss.